# A Unified Framework for 3D Scene Understanding

### v1

@@ 3.3 Explicit Inter-task Association
Associations among 3D scene understanding tasks have been overlooked in previous studies. It leads to a result that the task-specific approaches can not perceive across-task knowledge, limiting the understanding of the 3D scenes to a task-specific perspective. This limitation barries comprehensive 3D scene understanding. Thus, we build explicit inter-task associations to break this limitation.
Specifically, on the one hand, as shown in Fig. 3(a), the referring segmentation is challenging when multiple individuals of identical shapes are arranged adjacently. It requires the method to distinguish the location variations involved in the text prompts, such as ‘right of the other chair’ vs. ‘another chair to the right of it’. However, the modality gap between the point and text modalities set significant barriers. We propose ranking-based contrastive learning between the vision and text features to reduce the modality gap and optimize the referring segmentation.
Table 1: Mask prediction performance of interactive segmentation.
task mIoU
Interactive Seg. 76.0
On the other hand, as shown in Tab. 1, the mIoU metric measures the quality of mask predictions directly, we observe that the interactive segmentation task equips with superior mask prediction performance. We attribute it to the reliable position priors from the vision prompts. Consequently, we transfer knowledge from the interactive segmentation to the rest of the tasks to guide the training phase.
Figure 3: Illustration of the inter-task association. (a) A confusing case where requiring to distinguish the textual position information involved in the expressions. (b) A contrastive learning matrix for the vision-text pairs. The ranking rule is designed to suppress the incorrect pairs. (c) Knowledge distillations between multi-task predictions.
@@ 3.3.1 Ranking-based Contrastive Learning
We set the point and text prompts specifying the same individual instances into pairs and align their pair-wise features by employing contrastive learning.
Assuming $B$ text-point pairs within a training mini-batch, the corresponding output features are $\left\{\mathbf{f}_{out,i}^{p}\right\}_{i=1}^{B}$ and $\left\{\mathbf{f}_{out,i}^{t}\right\}_{i=1}^{B}$, where $\mathbf{f}_{out,i}^{p}$ and $\mathbf{f}_{out,i}^{t}$ are selected from output features $\left\{\mathbf{f}_{out,i}\right\}_{i=m+1}^{m+K_{p}}$ and $\left\{\mathbf{f}_{out,i}\right\}_{i=m+K_{p}+1}^{m+K_{p}+K_{t}}$, respectively. We normalize the pair-wise point-text output features $\left\{\mathbf{f}_{out,i}^{p}\right\}_{i=1}^{B}$ and $\left\{\mathbf{f}_{out,i}^{t}\right\}_{i=1}^{B}$ and obtain the metric embeddings $\left\{\mathbf{e}_{i}^{p}\right\}_{i=1}^{B}$ and $\left\{\mathbf{e}_{i}^{t}\right\}_{i=1}^{B}$, respectively. We formulate the contrastive learning loss as $\mathcal{L}_{con}=\mathcal{L}_{p}+\mathcal{L}_{t}$, with:
$\mathcal{L}_{p}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\left(\mathbf{e}_{i}^{p}\mathbf{e}_{i}^{t}/\tau\right)}{\sum_{j=1}^{B}\exp\left(\mathbf{e}_{i}^{p}\mathbf{e}_{j}^{t}/\tau\right)},~{}~{}~{}\mathcal{L}_{t}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\left(\mathbf{e}_{i}^{t}\mathbf{e}_{i}^{p}/\tau\right)}{\sum_{j=1}^{B}\exp\left(\mathbf{e}_{i}^{t}\mathbf{e}_{j}^{p}/\tau\right)},$ (4)
where $\tau$ is a learnable temperature. The pair-wise similarity can be depicted in Fig. 3(b), where we denote $\mathbf{e}_{i}^{p}\mathbf{e}_{j}^{t}$ as $\mathbf{s}_{i,j}$ for simplification. To distinguish the target instances from adjacent ones with identical shapes, we introduce a ranking rule inspired by the CrowdCLIP liang2023crowdclip that the diagonal elements are greater than the off-diagonal elements, which can be described as:
@@ 3.3.2 Knowledge Distillation
As shown in Fig. 3(c), we transfer knowledge from the interactive task to the generic and referring tasks to guide their training phases.
Interactive segmentation to the generic segmentation task. Define the predictions by the unified queries as $\mathbf{Pred}_{u}=\left\{\mathbf{mask}_{i},\mathbf{cls}_{i}\right\}_{i=1}^{m}$. We employ the Hungarian algorithm, utilizing the Dice and cross-entropy metrics as the matching cost criteria to assign $\mathbf{Pred}_{u}$ with the interactive segmentation labels $\mathbf{GT}_{p}=\left\{\mathbf{mask}_{i}^{gt},\mathbf{cls}_{i}^{gt}\right\}_{i=1}^{K_{p}}$, and obtain the positive samples as $\mathbf{Pos}_{u}=\left\{\mathbf{mask}_{i}^{pos},\mathbf{cls}_{i}^{pos}\right\}_{i=1}^{K_{p}}$. The predicted masks by the interactive segmentation can be formulated as $\mathbf{mask}_{p}=\left\{\mathbf{mask}_{i}\right\}_{i=m+1}^{m+K_{p}}\in\mathbb{R}^{K_{p}\times m}$. We select the pixels with top $k\%$ scores as learning region $\mathbf{R}\subseteq\mathbf{mask}_{p}$, and depict the knowledge transfer process from the interactive segmentation to the generic segmentation task as:
$\mathcal{L}_{p\to g}=\mathcal{L}_{BCE}\left(\mathbf{mask}_{pos}\left(\mathbf{R}\right),\mathbf{mask}_{p}\left(\mathbf{R}\right)\right),$ (6)
where $\mathbf{mask}_{pos}\left(\mathbf{R}\right)$ and $\mathbf{mask}_{p}\left(\mathbf{R}\right)$ represent the estimated mask values gathered in the region $\mathbf{R}$ from the positive samples and the interactive segmentation predictions, respectively.
Interactive segmentation to the referring segmentation task. Define the pair-wise class probabilities predicted by the point and text queries as $\mathbf{cls}_{p}$ and $\mathbf{cls}_{t}$, respectively. We formulate a knowledge transfer process from the interactive segmentation to the referring segmentation task as
$\mathcal{L}_{p\to r}=\mathcal{L}_{BCE}\left(\mathrm{Sigmoid}\left(\mathbf{cls}_{t}\right),\mathrm{Sigmoid}\left(\mathbf{cls}_{p}\right)\right).$ (7)
@@ 3.4 Training Objectives
Open-set pseudo mask labels. For the open-vocabulary tasks, the methods are trained on the close set. To enhance the segmentation performance on the open-set, we use SAM3D yang2023sam3d to generate pseudo segmentation masks with undetermined categories. Specifically, as discussed in Sec. 3.2, the positive samples are used to regress the close-set masks. For the negative samples, we match them with the pseudo segmentation masks and select the matched ones to regress the pseudo labels. Note that the SAM3D is an unsupervised method and uses no ground-truth annotations, which means there are no worries about label leakage. And this process is only employed during the training phase, which does not increase the inference cost.
Loss function. The training losses contain two components: (1) the basic losses, formulated as $\mathcal{L}_{base}=\mathcal{L}_{mask}+\mathcal{L}_{cls}$. $\mathcal{L}_{mask}$ stands for pixel-wise mask loss, which comprises of the BCE loss and the Dice loss. $\mathcal{L}_{cls}$ indicates the classification loss, where we use the cross-entropy loss. (2) the losses used to build inter-task associations, summarized as $\mathcal{L}_{inter}=\mathcal{L}_{p\to g}+\mathcal{L}_{p\to r}+\mathcal{L}_{con}+\mathcal{L}_{rank}$. The final loss function is $\mathcal{L}=\mathcal{L}_{base}+\lambda\mathcal{L}_{inter}$, where $\lambda$ is a balance weight, setting as $0.1$.
@@ 4 Experiments
Datasets. We evaluate the UniSeg3D on three benchmarks: ScanNet20 dai2017scannet , ScanNet200 rozenberszki2022lground , and ScanRefer chen2020scanrefer . ScanNet20 provides RGB-D images and 3D point clouds of $1,613$ scenes, including $18$ instance categories and $2$ semantic categories. ScanNet200 uses the same source data as the ScanNet20, while it is more challenging for up to $198$ instance categories and $2$ semantic categories. ScanRefer contains $51,583$ natural language expressions referring to $11,046$ objects selected from $800$ scenes.
Experimental setups. We train our method on the ScanNet20 training split, and the referring texts are collected from the ScanRefer. The $d_{in}$ and $d_{out}$ are set as $32$ and $256$, respectively. $m$ ranges $[50,100]$ percents of $M$ with an upper limit of $3,500$. We set $k$ as $10$ and $L$ as $6$. For the data augmentations, the point clouds are randomly flipped horizontally, rotated around the z-axis, elastic distorted, and scaled; the referring texts are augmented using public GPT tools following wu2023language ; dai2023auggpt . We adopt the AdamW optimizer with the polynomial schedule, setting an initial learning rate as $0.0001$ and the weight decay as $0.05$. All models are trained for $512$ epochs on a single NVIDIA RTX 4090 GPU and evaluated per $16$ epochs on the validation set to find the best-performed model. To stimulate the performance, we propose a two-stage fine-tuning trick, which fine-tunes the best-performed model using the learning rate and weight decay $0.001$ times the initial values for $40$ epochs. The proposed framework achieves end-to-end generic, interactive, and referring segmentation tasks. We divide the OVS task into mask prediction and class prediction. Specifically, we employ the proposed UniSeg3D to predict masks and then follow the Open3DIS nguyen2023open3dis to generate class predictions.
As for the metrics, we use the PQ, mIoU, and mAP to evaluate performance on the generic tasks following narita2019panopticfusion ; qi2017pointnet++ ; yi2019gspn . Then, we use the AP and mIoU for the interactive and referring segmentation tasks, respectively, following yue2023agile3d ; qian2024x . For the OVS task, we train our model on the ScanNet20 and evaluate it using the AP metric on the ScanNet200 without specific fine-tuning, following takmaz2024openmask3d .
@@ 4.1 Comparison to the SOTA Methods
The proposed method achieves six 3D scene understanding tasks in a single model. We demonstrate the effectiveness of our method by comparing it with SOTA approaches specialized for specific tasks. As shown in Tab. LABEL:tab:comparison, the proposed method outperforms the specialized SOTA methods PanopticNDT seichter2023panopticndt , OctFormer wang2023octformer , MAFT lai2023mask , AGILE3D yue2023agile3d , X-RefSeg3D qian2024x , and Open3DIS nguyen2023open3dis on the PS, SS, IS, interactive segmentation, referring segmentation, and OVS tasks by $12.1$ PQ, $1.2$ mIoU, $0.9$ mAP, $1.0$ AP, $4.1$ mIoU, $0.7$ AP, respectively. Even when compared with the competitive 3D unified method, i.e., OneFormer3D kolodiazhnyi2023oneformer3d , the proposed UniSeg3D achieves $0.1$ PQ improvement on the PS task, and $0.3$ mIoU improvement on the SS task. More importantly, the OneFormer3D focuses on three generic segmentation tasks. It fails to understand user prompts and is unable to achieve OVS, which limits its application prospects. In contrast, the UniSeg3D unifies six tasks and demonstrates desirable performance across all of them. This illustrates that the UniSeg3D is a powerful architecture.
Table 2: Ablation on components. ‘Distillation’, ‘Rank-Contrastive’, and ‘Trick’ denote the knowledge distillation, ranking-based contrastive learning, and two-stage fine-tuning trick, respectively.
Datasets ScanNet20 ScanRefer ScanNet200
Components PS SS IS Interactive Referring OVS
Distillation Rank-Contrastive Trick PQ mIoU mAP AP mIoU AP
- - - 70.4 76.2 58.0 54.5 29.1 19.7
- - 71.2 76.2 59.3 56.6 29.2 19.6
- - 70.8 76.4 58.4 54.1 29.6 20.2
- 71.3 76.3 59.1 54.1 29.5 19.6
71.3 76.9 59.3 54.5 29.6 19.7
Table 3: Ablation on different designs of the proposed components. ‘${p\to g}$’ and ‘${p\to r}$’ denote the knowledge distillation from the interactive segmentation to the generic segmentation and the referring segmentation, respectively. ‘Contrastive’ and ‘Rank’ denote the contrastive learning and the ranking rule, respectively.
(a) Ablation on designs for knowledge distillation.
${p\to g}$ ${p\to r}$ PS SS IS Interactive Referring OVS
- - 70.8 76.4 58.4 54.1 29.6 20.2
- 71.4 76.3 59.0 54.0 29.6 19.8
- 70.7 76.2 58.4 54.1 29.7 20.1
71.3 76.3 59.1 54.1 29.5 19.6
(b) Ablation on designs for ranking-based contrastive learning.
Contrastive Rank PS SS IS Interactive Referring OVS
- - 71.2 76.2 59.3 56.6 29.2 19.6
- 70.9 76.3 59.0 53.8 29.4 19.7
- 70.8 76.2 58.6 53.4 29.4 19.8
71.3 76.3 59.1 54.1 29.5 19.6
Figure 4: Ablation on hyper-parameter $\lambda$.
$\lambda$ PS SS IS Interactive Referring OVS
0.05 70.7 76.2 58.9 54.4 29.5 19.6
0.1 71.3 76.3 59.1 54.1 29.5 19.6
0.2 70.8 76.6 58.6 52.3 29.8 19.5
0.3 70.6 75.7 58.4 51.6 29.6 19.3
Figure 5: Ablation on task unification.
OVS Referring Interactive PS SS IS
@@
The proposed method achieves six tasks in once training, which is elegant while facing an issue for fair comparison that part of the labels ($10,115$ objects, $27.6$ percent of the complete ScanRefer training set) in the referring segmentation benchmark belongs to the novel classes of the open-vocabulary segmentation task. Obviously, these labels should not be used for training to avoid label leakage. Thus, we filter out this part of labels and only employ the filtered ScanRefer training set to train our model. As we can see in Tab. LABEL:tab:comparison, the proposed method use $72.4$ percent of the training data to achieve closing performance with X-RefSeg3D qian2024x ($29.6$ vs. $29.9$), the current specialized SOTA on the 3D referring segmentation task. Moreover, while we reproduce the X-RefSeg3D using official code on the same set of filtered training data, the performance drops to $4.1$ mIoU lower than ours. It demonstrates the effectiveness of our method on the referring segmentation task.

### v2

@@ 3.3 Explicit Inter-task Association
Previous studies have overlooked the associations among 3D scene understanding tasks, resulting in task-specific approaches that fail to leverage cross-task knowledge. This limitation restricts the understanding of 3D scenes to a task-specific perspective, hindering comprehensive 3D scene understanding. We establish explicit inter-task associations to overcome these constraints.
Specifically, on the one hand, as shown in Fig. 3(a), the referring segmentation is challenging when multiple individuals of identical shapes are arranged adjacently. It requires the method to distinguish the location variations inserted in the text prompts, such as “right of the other chair” vs. “another chair to the right of it.” However, the modality gap between 3D points and linguistic texts sets significant obstructions. We propose ranking-based contrastive learning between the vision and text features to reduce the modality gap and optimize the referring segmentation.
Table 1: Mask prediction performance of instance and interactive segmentation.
Tasks mIoU
Instance Seg. 68.1
Interactive Seg. 76.0 (+7.9)
On the other hand, as shown in Tab. 1, we evaluate our baseline framework built in Sec. 3.1 and Sec. 3.2 on instance and interactive segmentation tasks. Essentially, the main difference between the instance and interactive segmentation is w/o or w/ vision prompts. The mIoU metric, which directly measures the quality of mask predictions, indicates that the interactive segmentation surpasses the instance segmentation by a notable margin of $7.9\%$. It suggests that the vision prompts provide reliable position priors, boosting the interactive segmentation to perform superior mask prediction performance. We design a knowledge distillation to share insights from the interactive segmentation, leveraging its superior mask prediction capability. The key to knowledge distillation is to utilize task-predicting segmentation masks of the best quality to guide the other tasks, i.e., using a teacher to guide students.
@@ 3.3.1 Ranking-based Contrastive Learning
We set the vision and text prompts specifying the same individual instances into pairs and align their pair-wise features by employing contrastive learning.
Assuming $B$ vision-text pairs within a training mini-batch, the corresponding output features are $\left\{\mathbf{f}_{out,i}^{v}\right\}_{i=1}^{B}$ and $\left\{\mathbf{f}_{out,i}^{t}\right\}_{i=1}^{B}$, where $\mathbf{f}_{out,i}^{v}$ and $\mathbf{f}_{out,i}^{t}$ are selected from output features $\left\{\mathbf{f}_{out,i}\right\}_{i=m+1}^{m+K_{v}}$ and $\left\{\mathbf{f}_{out,i}\right\}_{i=m+K_{v}+1}^{m+K_{v}+K_{t}}$, respectively. We normalize the pair-wise vision-text output features $\left\{\mathbf{f}_{out,i}^{v}\right\}_{i=1}^{B}$ and $\left\{\mathbf{f}_{out,i}^{t}\right\}_{i=1}^{B}$ and obtain the metric embeddings $\left\{\mathbf{e}_{i}^{v}\right\}_{i=1}^{B}$ and $\left\{\mathbf{e}_{i}^{t}\right\}_{i=1}^{B}$, respectively. We formulate the contrastive learning loss as $\mathcal{L}_{con}=\mathcal{L}_{v}+\mathcal{L}_{t}$, with:
$\mathcal{L}_{v}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\left(\mathbf{e}_{i}^{v}\cdot\mathbf{e}_{i}^{t}/\tau\right)}{\sum_{j=1}^{B}\exp\left(\mathbf{e}_{i}^{v}\cdot\mathbf{e}_{j}^{t}/\tau\right)},~{}~{}~{}\mathcal{L}_{t}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\left(\mathbf{e}_{i}^{t}\cdot\mathbf{e}_{i}^{v}/\tau\right)}{\sum_{j=1}^{B}\exp\left(\mathbf{e}_{i}^{t}\cdot\mathbf{e}_{j}^{v}/\tau\right)},$ (4)
where $\tau$ is a learnable temperature scaling factor. The pair-wise similarity is illustrated in Fig. 3(b), where we denote $\mathbf{e}_{i}^{v}\cdot\mathbf{e}_{j}^{t}$ as $\mathbf{s}_{i,j}$ for simplification. To distinguish the target instances from adjacent ones with identical shapes, we introduce a ranking rule inspired by the CrowdCLIP liang2023crowdclip that the diagonal elements are greater than the off-diagonal elements, which can be described as:
@@ 3.3.2 Knowledge Distillation
As shown in Fig. 3(c), we transfer knowledge from the interactive segmentation task to the generic and referring segmentation tasks to guide their training phases.
Interactive segmentation to generic segmentation task. Define the predictions from the unified queries as $Pred_{u}=\left\{\mathbf{mask}_{i},\mathbf{cls}_{i}\right\}_{i=1}^{m}$. We employ the Hungarian algorithm, utilizing the Dice and cross-entropy metrics as the matching cost criteria, to assign $Pred_{u}$ with the interactive segmentation labels $GT_{v}=\left\{\mathbf{mask}_{gt,i},\mathbf{cls}_{gt,i}\right\}_{i=1}^{K_{v}}$. The matched predictions are selected as positive samples $Pos_{u}=\left\{\mathbf{mask}_{pos,i},\mathbf{cls}_{pos,i}\right\}_{i=1}^{K_{v}}$. The predicted masks from the interactive segmentation can be formulated as $\mathbf{mask}_{v}=\left\{\mathbf{mask}_{i}\right\}_{i=m+1}^{m+K_{v}}$, where $\mathbf{mask}_{v}\in\mathbb{R}^{K_{v}\times m}$. We select the pixels with top $k\%$ scores of $\mathbf{mask}_{v}$ as learning region $\mathbf{R}$, and depict the knowledge transfer process from the interactive segmentation to the generic segmentation task as:
$\mathcal{L}_{v\to g}=\mathcal{L}_{BCE}\left(\mathbf{mask}_{pos}\left(\mathbf{R}\right),\mathbf{mask}_{v}\left(\mathbf{R}\right)\right),$ (6)
where $\mathbf{mask}_{pos}\left(\mathbf{R}\right)$ and $\mathbf{mask}_{v}\left(\mathbf{R}\right)$ represent the predicted mask values within the region $\mathbf{R}$, gathering from the positive samples and the interactive segmentation predictions, respectively.
Interactive segmentation to referring segmentation task. Define the pair-wise class probabilities predicted by the vision and text prompt queries as $\mathbf{cls}_{v}\in\mathbb{R}^{B\times K_{c}}$ and $\mathbf{cls}_{t}\in\mathbb{R}^{B\times K_{c}}$ selected from $\left\{\mathbf{cls}_{i}\right\}_{i=m+1}^{m+K_{v}}$ and $\left\{\mathbf{cls}_{i}\right\}_{i=m+K_{v}+1}^{m+K_{v}+K_{t}}$, respectively. We formulate a knowledge transfer process from the interactive segmentation to the referring segmentation task as:
$\mathcal{L}_{v\to r}=\mathcal{L}_{BCE}\left(\mathrm{Sigmoid}\left(\mathbf{cls}_{t}\right),\mathrm{Sigmoid}\left(\mathbf{cls}_{v}\right)\right).$ (7)
@@ 3.4 Training Objectives
Open-set pseudo mask labels. For open-vocabulary tasks, we train models on close-set data. To enhance segmentation performance on open-set data, we use SAM3D yang2023sam3d to generate segmentation masks with undetermined categories as pseudo mask labels (open-set masks). While training, we assign predictions of the unified queries with ground-truth masks (close-set masks). The assigned and miss-assigned predictions are divided into positive and negative samples, respectively. The positive samples are supervised to regress the close-set masks. We match the negative samples with the pseudo mask labels and supervise the matched ones to regress the open-set masks. Note that the SAM3D is an unsupervised method and does not rely on ground-truth annotations, eliminating worries of label leakage. This process is exclusively applied in the training phase, incurring no extra inference cost.
Loss function. The training losses contain two components: (1) the basic losses, formulated as $\mathcal{L}_{base}=\mathcal{L}_{mask}+\mathcal{L}_{cls}$. $\mathcal{L}_{mask}$ stands for pixel-wise mask loss, which comprises of the BCE loss and the Dice loss. $\mathcal{L}_{cls}$ indicates the classification loss, where we use the cross-entropy loss. (2) the losses used to build inter-task associations, summarized as $\mathcal{L}_{inter}=\mathcal{L}_{v\to g}+\mathcal{L}_{v\to r}+\mathcal{L}_{con}+\mathcal{L}_{rank}$. The final loss function is $\mathcal{L}=\mathcal{L}_{base}+\lambda\mathcal{L}_{inter}$, where $\lambda$ is a balance weight, setting as $0.1$.
@@ 4 Experiments
Datasets. We evaluate the UniSeg3D on three benchmarks: ScanNet20 dai2017scannet , ScanNet200 rozenberszki2022lground , and ScanRefer chen2020scanrefer . ScanNet20 provides RGB-D images and 3D point clouds of $1,613$ scenes, including $18$ instance categories and $2$ semantic categories. ScanNet200 uses the same source data as ScanNet20, while it is more challenging for up to $198$ instance categories and $2$ semantic categories. ScanRefer contains $51,583$ natural language expressions referring to $11,046$ objects selected from $800$ scenes.
Experimental setups. We train our method on the ScanNet20 training split, and the referring texts are collected from the ScanRefer. The $d_{in}$ and $d_{out}$ are set as $32$ and $256$, respectively. $m$ ranges $[50,100]$ percents of $M$ with an upper limit of $3,500$. We set $k$ as $10$ and $L$ as $6$. For the data augmentations, the point clouds are randomly rotated around the z-axis, elastic distorted, and scaled; the referring texts are augmented using public GPT tools following wu2023language ; dai2023auggpt . We adopt the AdamW optimizer with the polynomial schedule, setting an initial learning rate as $0.0001$ and the weight decay as $0.05$. All models are trained for $512$ epochs on a single NVIDIA RTX 4090 GPU and evaluated per $16$ epochs on the validation set to find the best-performed model. To stimulate the performance, we propose a two-stage fine-tuning trick, which fine-tunes the best-performed model using the learning rate and weight decay $0.001$ times the initial values for $40$ epochs. The proposed framework achieves end-to-end generic, interactive, and referring segmentation tasks. We divide the OVS task into mask prediction and class prediction. Specifically, we employ the proposed UniSeg3D to predict masks and then follow the Open3DIS nguyen2023open3dis to generate class predictions.
We use PQ, mIoU, and mAP metrics to evaluate performance on the generic tasks following narita2019panopticfusion ; qi2017pointnet++ ; yi2019gspn . Then, we use AP metric and mIoU for the interactive and referring segmentation tasks, respectively, following yue2023agile3d ; qian2024x . For the OVS task, we train our model on the ScanNet20 and evaluate it using AP metric on the ScanNet200 without specific fine-tuning, following takmaz2024openmask3d . The Overall metric represents the average performance across six tasks intended to reflect the model’s unified capability.
@@ 4.1 Comparison to the SOTA Methods
The proposed method achieves six 3D scene understanding tasks in a single model. We demonstrate the effectiveness of our method by comparing it with SOTA approaches specialized for specific tasks. As shown in Tab. LABEL:tab:comparison, the proposed method outperforms the specialized SOTA methods PanopticNDT seichter2023panopticndt , OctFormer wang2023octformer , MAFT lai2023mask , AGILE3D yue2023agile3d , X-RefSeg3D qian2024x , and Open3DIS nguyen2023open3dis on the panoptic segmentation (PS), semantic segmentation (SS), instance segmentation (IS), interactive segmentation, referring segmentation, and OVS tasks by $12.1$ PQ, $1.2$ mIoU, $0.9$ mAP, $1.0$ AP, $4.1$ mIoU, $0.7$ AP, respectively. Even when compared with the competitive 3D unified method, i.e., OneFormer3D kolodiazhnyi2023oneformer3d , the proposed UniSeg3D achieves $0.1$ PQ improvement on the PS task, and $0.3$ mIoU improvement on the SS task. More importantly, the OneFormer3D focuses on three generic segmentation tasks. It fails to understand user prompts and achieve OVS, which limits its application prospects. In contrast, UniSeg3D unifies six tasks and presents desirable performance, demonstrating UniSeg3D a powerful architecture.
Table 2: Ablation on task unification.
ScanNet200 ScanRefer ScanNet20
OV Ref. Inter. Pan. Sem. Inst.
@@
The proposed method achieves six tasks in one training, which is elegant while facing an issue for fair comparison. Specifically, partial labels in the referring segmentation benchmark ($10,115$ objects, $27.6\%$ of the complete ScanRefer training set) annotate novel classes of the OVS task. Obviously, these labels should not be used for training to avoid label leakage. Thus, we filter out these labels and only employ the filtered ScanRefer training set to train our model. As shown in Tab. LABEL:tab:comparison, our model uses $72.4\%$ training data to achieve closing performance with X-RefSeg3D qian2024x ($29.6$ vs. $29.9$), the current specialized SOTA on the 3D referring segmentation task. Moreover, while reproducing the X-RefSeg3D using official code on our filtered training data, the performance drops to $4.1$ mIoU lower than UniSeg3D, demonstrating our model’s effectiveness.